assume: p(a)=.4,p(b)=.7 and p(AuB=.3. find p(AuB) a. .4 b..8c..2 d. .1

assume:

p(a)=.4,p(b)=.7 and p(AuB=.3.

find p(AuB)

a. .4 b..8c..2 d. .1
Answer #1

P(A∩B) = P(A) + P(B) - P(A ∪ B)

we have,
P(A) = 0.4
P(B) = 0.7
P(A U B) = 0.3
P(A∩B) = 0.4 + 0.7 - 0.3 = 0.8

please upvote if this helps